Assume that you have a capital of 100.

A friend suggests you to invest in his company in bonds X generating an interest of 10% for two years, while the bank is giving an interest of 5%per year. Can you determine what better investment is in terms of ROI? Put all your calculation on paper


1
Expert's answer
2021-04-29T10:27:06-0400

Case 1. An interest of 10% earns 100 x 0.1 = 10 dollars per year.

ROI = 100 / 10 = 10 years


Case 2. An interest rate of 5% earns 100 x 0.05 = 5 dollars per year

ROI = 100 / 5 = 20 years


Therefore, case 1 provides better conditions in terms of ROI.